# Diffusion Model as a Noise-Aware Latent Reward Model for Step-Level Preference Optimization

### v3

@@ Abstract
Preference optimization for diffusion models aims to align them with human preferences for images. Previous methods typically use Vision-Language Models (VLMs) as pixel-level reward models to approximate human preferences. However, when used for step-level preference optimization, these models face challenges in handling noisy images of different timesteps and require complex transformations into pixel space. In this work, we show that pre-trained diffusion models are naturally suited for step-level reward modeling in the noisy latent space, as they are explicitly designed to process latent images at various noise levels. Accordingly, we propose the Latent Reward Model (LRM), which repurposes components of the diffusion model to predict preferences of latent images at arbitrary timesteps. Building on LRM, we introduce Latent Preference Optimization (LPO), a step-level preference optimization method conducted directly in the noisy latent space. Experimental results indicate that LPO significantly improves the model’s alignment with general, aesthetic, and text-image alignment preferences, while achieving a 2.5-28$\times$ training speedup over existing preference optimization methods. Our code and models are available at https://github.com/Kwai-Kolors/LPO.
@@ 1 Introduction
footnotetext: 🖂Corresponding author: Kun Ding $<$kun.ding@ia.ac.cn$>$.
@@
Based on the above insights, we propose using the diffusion model as a noise-aware Latent Reward Model (LRM). Given noisy latent images $x_{t}$, LRM utilizes visual features from U-Net [12] or DiT [13] and textual features from the text encoder to predict step-level preference labels. A Visual Feature Enhancement (VFE) module is employed to enhance LRM’s focus on the text-image alignment. To address the inconsistent preference issue in LRM’s training data, we propose the Multi-Preference Consistent Filtering (MPCF) strategy to ensure that winning images consistently outperform losing ones across multiple dimensions. Finally, we employ LRM for step-level preference optimization, leading to a simple yet effective method termed Latent Preference Optimization (LPO), where all steps are conducted within the noisy latent space of diffusion models. Extensive experiments on SD1.5 [14] and SDXL [11] demonstrate that LPO substantially improves the quality of generated images and consistently outperforms existing DPO and SPO methods across the general, aesthetic, and alignment preferences, as indicated in Fig. 1. Meanwhile, LPO exhibits remarkable training efficiency, achieving a speedup of 10-28$\times$ over Diffuison-DPO [4] and 2.5-3.5$\times$ over SPO [8]. Furthermore, we explore a step-wise variant of GRPO [15] based on LRM and apply LPO to the DiT-based SD3 [1] model, demonstrating the generalization ability of LRM and LPO.
@@ 3.1 Latent Diffusion Models
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{t},\sigma_{t}^{2}\epsilon_{t}),\quad\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\$ (2)
$\mu_{t}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0,t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta,t},\quad\hat{x}_{0,t}=(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta,t}}{\sqrt{\bar{\alpha}_{t}}}),\$ (3)
@@ 4.1 Latent Reward Model
Architecture of LRM. LRM leverages features of the U-Net and text encoder in the diffusion model for preference prediction, as depicted in Fig. 3. Specifically, the textual features $f_{p}\in\mathbb{R}^{n_{l}\times n_{p}}$ are extracted from the prompt $p$ by the text encoder, where $n_{l}$ and $n_{p}$ denote the number of tokens and the dimension of textual features, respectively. Following CLIP [9], we use the last token $f_{eos}\in\mathbb{R}^{1\times n_{p}}$ to represent the entire prompt and incorporate a text projection layer to obtain the final textual features $T\in\mathbb{R}^{1\times n_{d}}$, where $n_{d}$ represents the final dimension of textual and visual features. Each noisy latent image $x_{t}$ is passed through the U-Net to interact with textual features $f_{p}$. The visual features of U-Net are averagely pooled along the spatial dimension, resulting in the multiscale down-block features $V_{down}$ and middle-block features $V_{mid}$ as follows
@@
$L_{LRM}=-\mathbb{E}_{t\sim\mathcal{U}(0,T),(I^{w},I^{l},p)\in D}\log\frac{exp(S(p,x_{t}^{w}))}{exp(S(p,x_{t}^{w}))+exp(S(p,x_{t}^{l}))},\$ (10)
@@ 4.2 Multi-Preference Consistent Filtering
Issue of Inconsistent Preferences. The training loss of LRM in Eqn. (10) involves the assumption that if $x^{w}_{0}$ is preferred over $x^{l}_{0}$, then after adding noise of equal intensity, the preference order remains unchanged, meaning $x^{w}_{t}$ continues to be preferred over $x^{l}_{t}$. However, this assumption breaks down when the winning image excels in one aspect but is inferior in other aspects. For example, if the winning image $x^{w}_{0}$ has better details but exhibits weaker text-image alignment compared to the losing image $x_{0}^{l}$, the advantage in detail may be diminished after introducing significant noise, making the noisy losing image $x_{t}^{l}$ potentially more preferred than $x_{t}^{w}$. Such situations frequently occur in the Pick-a-Pic v1 [16] dataset. To demonstrate this, we employ the Aesthetic Score $S_{A}$ [38] to evaluate the aesthetic quality while using CLIP Score $S_{C}$ [9] and VQAScore $S_{V}$ [39] to assess the text-image alignment. Let $G_{*}=S_{*}(I^{w})-S_{*}(I^{l}),*\in\{A,C,V\}$ denote the preference score gaps between winning and losing images. As illustrated in Fig. 5, nearly half of the winning images have Aesthetic Scores lower than those of losing images. Similarly, around 40% of the winning images exhibit lower CLIP Scores and VQAScores. When considering both Aesthetic and CLIP scores, over 70% of the winning images score lower than the losing images in at least one aspect.
@@ 4.3 Latent Preference Optimization
To utilize LRM for step-level preference optimization, we propose the Latent Preference Optimization (LPO) method. We also explore the application of LRM on GRPO [15] in Appendix B.
@@
Optimization Timesteps. Due to the inaccuracy of SPM at high-noise levels, SPO samples training data only at low to medium noise levels, specifically when $t\in[0,750]$, as shown in Fig. 4 (c). In contrast, owing to LRM’s noise-aware ability, LPO can perform sampling and training throughout the entire denoising process, covering $t\in[0,950]$. Experiments in Sec. 5.3 and Appendix B demonstrate that optimization at high noise levels plays a crucial role in preference optimization.
@@ 5.1 Experimental Setup
LRM and LPO Training. We denote the LRM based on SD1.5 and SDXL as LRM-1.5 and LRM-XL, respectively. They are trained on the filtered Pick-a-Pic v1 [16] as clarified in Sec. 4.2. The $gs$ in the VFE module is set to 7.5. The same 4k prompts in SPO are used for the LPO training, randomly sampled from the training set of Pick-a-Pic v1. The DDIM scheduler [35] with 20 inference steps is employed. We use all steps for sampling and training, i.e. $t\in[0,50,...,900,950]$. The dynamic threshold range $[th_{min},th_{max}]$ is set to $[0.35,0.5]$ for SD1.5 and $[0.45,0.6]$ for SDXL. The $\beta$ in Eqn. (6) is set to 500 and the $K$ in the sampling process is set to 4. More details are in Appendix A.
Baseline Methods. We compare LPO with DDPO [6], D3PO [7], Diffusion-DPO [4], MaPO [29], SPO [8], and SePPO [30]. These methods are trained on Pick-a-Pic, which ensures a relatively fair comparison. We also include InterComp [31] for reference, which uses a higher-quality internal dataset. More details are provided in Appendix A.
Evaluation Protocol. We evaluate various diffusion models across three dimensions: general preference, aesthetic preference, and text-image alignment. The PickScore [16], HPSv2 [17], HPSv2.1 [17], and ImageReward [18] are utilized to assess the general preference. The aesthetic preference is evaluated using the Aesthetic Score [38]. Following [8], both general and aesthetic preferences are assessed on the validation unique split of Pick-a-Pic v1, which has 500 different prompts. For text-image alignment, we employ the GenEval [40] and T2I-CompBench++ [41] metrics. For detailed evaluation practice, please refer to Appendix A. Additionally, we propose two metrics to assess LRM’s correlations with aesthetics and text-image alignment. Specifically, we calculate the score gaps $G_{*},*\in\{A,C,L\}$ between winning and losing images, where $A$, $C$, $L$ represent Aesthetic, CLIP, and LRM. For LRM, the score is taken at $t=0$. Then the Pearson Correlation Coefficient [42] between $G_{L}$ and $G_{A}$ is referred to as Aes-Corr while that between $G_{L}$ and $G_{C}$ is termed CLIP-Corr. They are evaluated on the validation unique and test unique splits of Pick-a-Pic v1.
@@ 5.2 Main Results
Quantitative Comparison. As indicated in Tab. 2 and Tab. 3, LPO outperforms other methods across various dimensions by a large margin on SD1.5, especially in general preference and text-image alignment. Based on SDXL, LPO even slightly surpasses InterComp, although the latter utilizes a superior dataset comprising images from SD3[1] and FLUX [43]. The user study results in Appendix B further confirm the superiority of LPO. We also validate the effectiveness of heterogeneous optimization in Tab. 6. Remarkably, fine-tuning SD2.1 using LRM-1.5 yields significant improvements across various aspects, demonstrating that an inferior diffusion model can effectively fine-tune an advanced model as long as they share the same VAE encoder. In contrast, applying LRM-1.5 for the LPO of SDXL is ineffective due to a distribution mismatch between the latent spaces of their VAE encoders.
@@
Further Explorations. In Appendix B, we present additional exploration experiments on LRM and LPO. First, we validate the effectiveness of LRM on a step-wise variant of GRPO [15]. We then assess the applicability of LPO on DiT-Based SD3 [1] and evaluate its generalization on other datasets. Additionally, we report the performance of LRM when used as a pixel-wise reward model.
@@ 5.3 Ablation Studies
MPCF. As shown in Tab. 8, MPCF plays a vital role in LRM training. The first filtering strategy enforces that winning images must outperform losing images across all aspects. However, since the diffusion model lacks explicit text-image alignment pre-training like CLIP, it is prone to overfitting to visual features, as indicated by a higher Aes-Corr. This overfitting results in reduced attention to alignment, as reflected by lower CLIP-Corr and GenEval scores. The second and third strategies relax the aesthetic constraints to varying degrees. While the third, most lenient strategy can cause LRM to focus solely on alignment and neglect image quality, as evidenced by a negative Aes-Corr value, the second strategy achieves a better balance, yielding the highest general preference scores. Notably, even without MPCF and using the same training data, LPO still outperforms SPO in both general and text-image alignment preferences, as detailed in Appendix C.
@@
Optimization Timesteps. Tab. 10 ablates different optimization timestep ranges, indicating that broader ranges lead to better performance. Notably, LRM can effectively predict preferences even at very large timesteps, e.g., $[750,950]$, achieving results comparable to those obtained across the entire denoising process, i.e. $[0,950]$. This highlights the critical role of large timesteps in step-level preference optimization. In contrast, SPO performs poorly in large timesteps, as indicated in Tab. 23 in Appendix C. This comparison underscores the superiority of using diffusion models for step-level reward modeling at large timesteps, since LRM can directly process noisy latent images without suffering from distribution shift. A qualitative comparison of different ranges is provided in Fig. 10.
Dynamic Sampling Threshold. The standard deviation $\sigma_{t}$ of samples at smaller timesteps is relatively small according to the DDPM scheduling [44], making the constant threshold insufficient to accommodate all timesteps. As indicated in Tab. 10, the dynamic threshold strategy generally outperforms the constant threshold across different intervals, effectively alleviating this problem. We further explore other dynamic strategies in Tab. 26 in Appendix C.
@@ 6 Discussion and Conclusion
Limitations and Future Work. (1) The Pick-a-Pic dataset mainly contains images generated by SD1.5 and SDXL, which generally exhibit low image quality. Introducing higher-quality images is expected to enhance the generalization of the LRM. (2) As a step-level reward model, the LRM can be easily applied to reward fine-tuning methods [45, 24, 26], avoiding lengthy inference chain backpropagation and significantly accelerating the training speed. (3) The LRM can also extend the best-of-N approach to a step-level version, enabling exploration and selection at each step of image generation, thereby achieving inference-time optimization similar to GPT-o1 [46].
@@ Appendix A Experimantal Details
Training Resolution 512$\times$512 1024$\times$1024
@@ Appendix B Additional Exploration Experiments
We conduct preliminary experiments of step-wise GRPO on the SD1.5 model. We set $K$ to 4, $\beta$ to 0.1 and $\epsilon$ to 0.1. As shown in Tab. 14, based on our LRM-1.5, step-wise GRPO consistently improves the model’s performance across various dimensions. However, in the current setting, step-wise GRPO performs slightly worse than LPO and requires 8$\times$ the training time. We will further explore the application of LRM-based GRPO for image-domain preference optimization in future work.
@@ Appendix C Additional Ablation Experiments
Table 23: Ablataion results on the hyperparamter $\beta$.

### v5

@@ Abstract
[…]
@@ 1 Introduction
footnotetext: 🖂Corresponding author: Kun Ding <<kun.ding@ia.ac.cn>>.
@@
[…]
The core contributions of this paper are summarized as follows: (1) A noise-aware Latent Reward Model is introduced, which repurposes the pre-trained diffusion model for step-level reward modeling in the noisy latent space. (2) A Multi-Preference Consistent Filtering strategy is proposed to refine the public preference dataset, enabling LRM to better align with human preferences. (3) A Latent Preference Optimization method based on LRM is introduced to perform step-level preference optimization directly within the noisy latent space of diffusion models. (4) Extensive experimental results demonstrate the effectiveness, efficiency, and generalization ability of the proposed methods.
@@ 3.1 Latent Diffusion Models
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{t},\sigma_{t}^{2}\epsilon_{t}),\quad\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\\$ (2)
$\mu_{t}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0,t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta,t},\quad\hat{x}_{0,t}=(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta,t}}{\sqrt{\bar{\alpha}_{t}}}),\\$ (3)
@@ 4.1 Latent Reward Model
Architecture of LRM. LRM leverages features of the U-Net and text encoder in the diffusion model for preference prediction, as depicted in Fig. 3. Specifically, the textual features $f_{p}\in\mathbb{R}^{n_{l}\times n_{p}}$ are extracted from the prompt $p$ by the text encoder, where $n_{l}$ and $n_{p}$ denote the number of tokens and the dimension of textual features, respectively. Following CLIP [9], we use the last token $f_{eos}\in\mathbb{R}^{1\times n_{p}}$ to represent the entire prompt and incorporate a text projection layer to obtain the final textual features $T\in\mathbb{R}^{1\times n_{d}}$, where $n_{d}$ represents the final dimension of textual and visual features. Each noisy latent image $x_{t}$ is passed through the U-Net to interact with textual features $f_{p}$. The visual features of U-Net are average pooled along the spatial dimension, resulting in the multiscale down-block features $V_{down}$ and middle-block features $V_{mid}$ as follows
@@
$L_{LRM}=-\mathbb{E}_{t\sim\mathcal{U}(0,T),(I^{w},I^{l},p)\in D}\log\frac{exp(S(p,x_{t}^{w}))}{exp(S(p,x_{t}^{w}))+exp(S(p,x_{t}^{l}))},\\$ (10)
@@ 4.2 Multi-Preference Consistent Filtering
Issue of Inconsistent Preferences. The training loss of LRM in Eqn. (10) involves the assumption that if $x^{w}_{0}$ is preferred over $x^{l}_{0}$, then after adding noise of equal intensity, the preference order remains unchanged, meaning $x^{w}_{t}$ continues to be preferred over $x^{l}_{t}$. However, this assumption breaks down
@@
when the winning image excels in one aspect but is inferior in other aspects. For example, if the winning image $x^{w}_{0}$ has better details but exhibits weaker text-image alignment compared to the losing image $x_{0}^{l}$, the advantage in detail may be diminished after introducing significant noise, making the noisy losing image $x_{t}^{l}$ potentially more preferred than $x_{t}^{w}$. Such situations frequently occur in the Pick-a-Pic v1 [16] dataset. To demonstrate this, we employ the Aesthetic Score $S_{A}$ [38] to evaluate the aesthetic quality while using CLIP Score $S_{C}$ [9] and VQAScore $S_{V}$ [39] to assess the text-image alignment. Let $G_{*}=S_{*}(I^{w})-S_{*}(I^{l}),*\in\{A,C,V\}$ denote the preference score gaps between winning and losing images. As illustrated in Fig. 5, nearly half of the winning images have Aesthetic Scores lower than those of losing images. Similarly, around 40% of the winning images exhibit lower CLIP Scores and VQAScores. When considering both Aesthetic and CLIP scores, over 70% of the winning images score lower than the losing images in at least one aspect.
@@ 4.3 Latent Preference Optimization
To utilize LRM for step-level preference optimization, we propose the Latent Preference Optimization (LPO) method. We also explore the application of LRM on GRPO [15] in Appendix˜B.
@@
Optimization Timesteps. Due to the inaccuracy of SPM at high-noise levels, SPO samples training data only at low to medium noise levels, specifically when $t\in[0,750]$, as shown in Fig. 4 (c). In contrast, owing to LRM’s noise-aware ability, LPO can perform sampling and training throughout the entire denoising process, covering $t\in[0,950]$. Experiments in Sec. 5.3 and Appendix˜B demonstrate that optimization at high noise levels plays a crucial role in preference optimization.
@@ 5.1 Experimental Setup
LRM and LPO Training. We denote the LRM based on SD1.5 and SDXL as LRM-1.5 and LRM-XL, respectively. They are trained on the filtered Pick-a-Pic v1 [16] as clarified in Sec. 4.2. The $gs$ in the VFE module is set to 7.5. The same 4k prompts in SPO are used for the LPO training, randomly sampled from the training set of Pick-a-Pic v1. The DDIM scheduler [35] with 20 inference steps is employed. We use all steps for sampling and training, i.e. $t\in[0,50,...,900,950]$. The dynamic threshold range $th_{min},th_{max}]$ is set to $[0.35,0.5]$ for SD1.5 and $[0.45,0.6]$ for SDXL. The $\beta$ in Eqn. (6) is set to 500 and the $K$ in the sampling process is set to 4. More details are in Appendix˜A.
Baseline Methods. We compare LPO with DDPO [6], D3PO [7], Diffusion-DPO [4], MaPO [29], SPO [8], and SePPO [30]. These methods are trained on Pick-a-Pic, which ensures a relatively fair comparison. We also include InterComp [31] for reference, which uses a higher-quality internal dataset. More details are provided in Appendix˜A.
Evaluation Protocol. We evaluate various diffusion models across three dimensions: general preference, aesthetic preference, and text-image alignment. The PickScore [16], HPSv2 [17], HPSv2.1 [17], and ImageReward [18] are utilized to assess the general preference. The aesthetic preference is evaluated using the Aesthetic Score [38]. Following [8], both general and aesthetic preferences are assessed on the validation unique split of Pick-a-Pic v1, which has 500 different prompts. For text-image alignment, we employ the GenEval [40] and T2I-CompBench++ [41] metrics. For detailed evaluation practice, please refer to Appendix˜A. Additionally, we propose two metrics to assess LRM’s correlations with aesthetics and text-image alignment. Specifically, we calculate the score gaps $G_{*},*\in\{A,C,L\}$ between winning and losing images, where $A$, $C$, $L$ represent Aesthetic, CLIP, and LRM. For LRM, the score is taken at $t=0$. Then the Pearson Correlation Coefficient [42] between $G_{L}$ and $G_{A}$ is referred to as Aes-Corr while that between $G_{L}$ and $G_{C}$ is termed CLIP-Corr. They are evaluated on the validation unique and test unique splits of Pick-a-Pic v1.
@@ 5.2 Main Results
Quantitative Comparison. As indicated in Tab. 2 and Tab. 3, LPO outperforms other methods across various dimensions by a large margin on SD1.5, especially in general preference and text-image alignment. Based on SDXL, LPO even slightly surpasses InterComp, although the latter utilizes a superior dataset comprising images from SD3[1] and FLUX [43]. The user study results in Appendix˜B further confirm the superiority of LPO. We also validate the effectiveness of heterogeneous optimization in Tab. 6. Remarkably, fine-tuning SD2.1 using LRM-1.5 yields significant improvements across various aspects, demonstrating that an inferior diffusion model can effectively fine-tune an advanced model as long as they share the same VAE encoder. In contrast, applying LRM-1.5 for the LPO of SDXL is ineffective due to a distribution mismatch between the latent spaces of their VAE encoders.
@@
Further Explorations. In Appendix˜B, we present additional exploration experiments on LRM and LPO. First, we validate the effectiveness of LRM on a step-wise variant of GRPO [15]. We then assess the applicability of LPO on DiT-Based SD3 [1] and evaluate its generalization on other datasets. Additionally, we report the performance of LRM when used as a pixel-wise reward model.
@@ 5.3 Ablation Studies
MPCF. As shown in Tab. 8, MPCF plays a vital role in LRM training. The first filtering strategy enforces that winning images must outperform losing images across all aspects. However, since the diffusion model lacks explicit text-image alignment pre-training like CLIP, it is prone to overfitting to visual features, as indicated by a higher Aes-Corr. This overfitting results in reduced attention to alignment, as reflected by lower CLIP-Corr and GenEval scores. The second and third strategies relax the aesthetic constraints to varying degrees. While the third, most lenient strategy can cause LRM to focus solely on alignment and neglect image quality, as evidenced by a negative Aes-Corr value, the second strategy achieves a better balance, yielding the highest general preference scores. Notably, even without MPCF and using the same training data, LPO still outperforms SPO in both general and text-image alignment preferences, as detailed in Appendix˜C.
@@
Optimization Timesteps. Tab. 10 ablates different optimization timestep ranges, indicating that broader ranges lead to better performance. Notably, LRM can effectively predict preferences even at very large timesteps, e.g., $[750,950]$, achieving results comparable to those obtained across the entire denoising process, i.e. $[0,950]$. This highlights the critical role of large timesteps in step-level preference optimization. In contrast, SPO performs poorly in large timesteps, as indicated in Tab. 23 in Appendix˜C. This comparison underscores the superiority of using diffusion models for step-level reward modeling at large timesteps, since LRM can directly process noisy latent images without suffering from distribution shift. A qualitative comparison of different ranges is provided in Fig. 10.
Dynamic Sampling Threshold. The standard deviation $\sigma_{t}$ of samples at smaller timesteps is relatively small according to the DDPM scheduling [44], making the constant threshold insufficient to accommodate all timesteps. As indicated in Tab. 10, the dynamic threshold strategy generally outperforms the constant threshold across different intervals, effectively alleviating this problem. We further explore other dynamic strategies in Tab. 26 in Appendix˜C.
@@ 6 Discussion and Conclusion
Limitations. (1) The Pick-a-Pic dataset mainly contains images generated by SD1.5 and SDXL, which generally exhibit low image quality. Introducing higher-quality images is expected to enhance the generalization of the LRM. (2) MPCF relies on three reward models to approximate human preferences for automatic data filtering. However, the filtered data may still inherit biases or limitations shared by these reward models. (3) Since LPO is performed within the latent space, which is determined by the VAE encoder, the VAE of LRM should be identical to that of DMO.
Future Work. (1) As a step-level reward model, the LRM can be easily applied to reward fine-tuning methods [45, 24, 26], avoiding lengthy inference chain backpropagation and significantly accelerating the training speed. (2) The LRM can also extend the best-of-N approach to a step-level version, enabling exploration and selection at each step of image generation, thereby achieving inference-time optimization similar to GPT-o1 [46].
@@ Appendix A Experimantal Details
Training Resolution $\times512$ $\times1024$
@@ Appendix B Additional Exploration Experiments
[…]
@@ Appendix C Additional Ablation Experiments
Table 23: Ablation results on the hyperparamter $\beta$.
@@ Appendix D Additional Visualizations
Appendix E Additional Discussions
During the research process, we observed the following types of failure cases:
  • Reward hacking: When training steps are excessive, the model exhibits reward hacking, where the reward metric continues to improve while the quality of generated images degrades. This stems from a certain misalignment between the reward model and human preferences, which remains a common issue across most existing methods. Mitigating this problem requires a combination of strategies, including improvements in preference data and reward modeling approaches.
  • Decline in diversity: Extended training also leads to a reduction in generation diversity. This is caused by the reward model narrowing the output distribution of the diffusion model. A general solution is prompt engineering, such as rephrasing input prompts to enhance the diversity of generated images via prompt variation.
  • Lack of fine details on SDXL: For some prompts, SDXL models optimized with LPO generate images with less detailed content compared to those optimized with SPO. This phenomenon was not observed on SD1.5. We will investigate this in future work to identify the underlying causes.
Appendix F Broader Impacts